# Error Bound Analysis for the Regularized Loss of Deep Linear Neural Networks

### v1

@@ B.2 The Case $L\geq 3$
Applying the Taylor expansion to $f(\sigma^{*}_{i}+t)$ at $\sigma^{*}$, together with $f(\sigma^{*}_{i})=0$ and $f^{\prime}(\sigma^{*}_{i})=0$, yields that when $t\to 0$, $\|\nabla G(\bm{W}(t))\|_{F}=O(t^{2}).$ We also note that
@@
$\mathrm{dist}^{2}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})=\|\bm{W}_{1}(t)-\bm{W}_{1}^{*}\|_{F}^{2}+\cdots+\|\bm{W}_{L}(t)-\bm{W}_{L}^{*}\|_{F}^{2}\geq Lt^{2},$
where the inequality follows from Weyl’s inequality. Using the same argument in Lemma 6, we conclude that $\mathrm{dist}(\bm{W}(t),\widehat{\mathcal{W}}_{\bm{\sigma}^{*}})=\mathrm{dist}(\bm{W}(t),\mathcal{W}_{G})$ when $t$ is sufficient small. Then we have $\|\nabla G(\bm{W})\|_{F}=O\left(\mathrm{dist}^{2}(\bm{W}(t),\mathcal{W}_{G})\right)$, which implies that the error bound fails to hold. ∎
Remark 3.
Under 1, it follows from 2 that 2 is a sufficient condition for the error bound to hold for the critical point set $\mathcal{W}_{G}$. According to Lemma 6 and Lemma 8, we conclude that 2 is also a necessary condition. Therefore, we establish the sufficient and necessary condition for the error bound of the critical point set $\mathcal{W}_{G}$.
@@ Appendix C Linear Convergence under the Error Bound
    (Sufficient Decrease) There exists a constant $\kappa_{1}>0$ such that
@@
    (Cost-to-Go Estimate) There exists a constant $\kappa_{2}>0$ such that
@@
    (Safeguard) There exists a constant $\kappa_{3}>0$ such that
    $\|\nabla F(\bm{W}^{k})\|_{F}\leq\kappa_{3}\|\bm{W}^{k+1}-\bm{W}^{k}\|_{F}.$
Then, the sequence $\{F(\bm{W}^{k})\}_{k\geq k_{1}}$ converges Q-linearly to $F(\bm{W}^{*})$ and $\{\bm{W}^{k}\}_{k\geq k_{1}}$ converges R-linearly to some $\bm{W}^{*}\in\mathcal{W}$.
We should point out that first-order methods, such as GD and proximal alternating minimization method, for solving Problem (2) satisfy the above conditions provided that an appropriate step size is chosen.
C.1 Proof of the PŁ Inequality and Quadratic Growth
Lemma 9 (Local Lipschitz Property).
Let $\bm{W}$ and $\bar{\bm{W}}$ be arbitrary such that
$\mathrm{dist}(\bm{W},\mathcal{W})\leq\delta_{3}\ \text{and}\ \mathrm{dist}(\bar{\bm{W}},\mathcal{W})\leq\delta_{3},$ (189)
where $\delta_{3}>0$ is a constant and $\mathcal{W}$ denotes the critical point set. Then, it holds that
$\|\nabla F(\bm{W})-\nabla F(\bar{\bm{W}})\|_{F}\leq L_{F}\|\bm{W}-\bar{\bm{W}}\|_{F},$
where $L_{F}>0$ is a constant.
Proof.
According to 1, we note that $\mathcal{\bm{W}}$ is a compact set. It implies that there exists a positve constant $M$ such that $\|\bm{W}^{*}_{l}\|\leq M$ for each $l\in[L]$. This, together with (189) and the triangular inequality yields that there exists a constant $M$ such that
$\|\bm{W}_{l}\|\leq M+\delta_{3},\ \|\bar{\bm{W}}_{l}\|\leq M+\delta_{3},\ \forall l\in[L].$
Then, we have
$\frac{1}{2}\nabla_{\bm{W}_{l}}F(\bm{W})-\frac{1}{2}\nabla_{\bar{\bm{W}}_{l}}F(\bar{\bm{W}})=\lambda_{l}(\bm{W}_{l}-\bar{\bm{W}}_{l})+\bm{R}_{1}+\bm{R}_{2},$
where
$\bm{R}_{1}$ $:=(\bm{W}_{l+1}-\bar{\bm{W}}_{l+1})^{T}\bm{W}_{L:l+2}^{T}\bm{W}_{L:1}\bm{W}_{l-1:1}^{T}+\cdots+\bar{\bm{W}}_{L:l+1}^{T}\bar{\bm{W}}_{L:1}\bar{\bm{W}}_{l-2:1}^{T}(\bm{W}_{l-1}-\bar{\bm{W}}_{l-1})^{T},$
$\bm{R}_{2}$ $:=(\bm{W}_{l+1}-\bar{\bm{W}}_{l+1})^{T}\bm{W}_{L:l+2}^{T}\bm{Y}\bm{W}_{l-1:1}^{T}+\cdots+\bar{\bm{W}}_{L:l+1}^{T}\bm{Y}\bar{\bm{W}}_{l-2:1}^{T}(\bm{W}_{l-1}-\bar{\bm{W}}_{l-1})^{T}.$
Then we compute
$\|\bm{R}_{1}\|_{F}+\|\bm{R}_{2}\|_{F}\leq\left(\|\bm{Y}\|(L-1)(M+\delta_{3})^{L-1}+(2L-1)(M+\delta_{3})^{2L-1}\right)\|\bm{W}_{l}-\bar{\bm{W}}_{l}\|_{F}.$
It implies that
$\|\nabla F(\bm{W})-\nabla F(\bar{\bm{W}})\|_{F}\leq 2L(\lambda_{\max}+L(M+\delta_{3})^{L-1}\|\bm{Y}\|+2L(M+\delta_{3})^{2L-1})\|\bm{W}-\bar{\bm{W}}\|_{F}.$
Proof of 1.
(i) Let $\mathcal{U}$ be a neighborhood in which both 2 and Lemma 9 hold and $F(\bm{W})\leq F(\bm{W}^{*})+\nu$, where $\nu>0$ is a constant. For any $\bm{W}\in\mathcal{U}$, let $\bm{W}^{*}\in\mathcal{W}$ be such that
$\mathrm{dist}(\bm{W},\mathcal{W})=\|\bm{W}-\bm{W}^{*}\|_{F}.$ (190)
According to Lemma 9, it follows that
$F(\bm{W})$ $\leq F(\bm{W}^{*})+\langle\nabla F(\bm{W}^{*}),\bm{W}-\bm{W}^{*}\rangle+\frac{L_{F}}{2}\|\bm{W}-\bm{W}^{*}\|_{F}^{2}$
$=F(\bm{W}^{*})+\frac{L_{F}}{2}\|\bm{W}-\bm{W}^{*}\|_{F}^{2}\leq F(\bm{W}^{*})+\frac{\kappa_{1}L_{F}}{2}\|\nabla F(\bm{W})\|_{F}^{2},$
where the equality is because $\bm{W}^{*}$ is a critical point, and the last inequality uses (11) and (190).
(ii) This result has been established in [34, Proposition 2.2]. ∎
@@ Appendix D Auxiliary Results
Lemma 10.
Let $\bm{A},\hat{\bm{A}}\in\mathbb{R}^{p\times p}$ be symmetric with eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{p}$ and $\hat{\lambda}_{1}\geq\cdots\geq\hat{\lambda}_{p}$ respectively. Fix $1\leq r\leq s\leq p$ and assume that $\min\{\lambda_{r-1}-\lambda_{r},\lambda_{s}-\lambda_{s+1}\}>0$, where $\lambda_{0}:=\infty$ and $\lambda_{p+1}:=-\infty$. Let $d:=s-r+1$ and $\bm{V}=(\bm{v}_{r},\bm{v}_{r+1},\dots,\bm{v}_{s})\in\mathbb{R}^{p\times d}$ and $\hat{\bm{V}}=(\hat{\bm{v}}_{r},\hat{\bm{v}}_{r+1},\dots,\hat{\bm{v}}_{s})\in\mathbb{R}^{p\times d}$ have orthonormal columns satisfying $\bm{A}\bm{v}_{j}=\lambda_{j}\bm{v}_{j}$ and $\hat{\bm{A}}\hat{\bm{v}}_{j}=\hat{\lambda}_{j}\hat{\bm{v}}_{j}$ for all $j=r,r+1,\ldots,s$. Then, there exists an orthogonal matrix $\bm{O}\in\mathcal{O}^{d}$ such that
@@
Lemma 11 (Mirsky Inequality [37]).
For any matrices $\bm{X},\tilde{\bm{X}}\in\mathbb{R}^{m\times n}$ with singular values
@@
where $l=\min\{m,n\}$, then for any unitarily invariant norm (e.g., $\|\cdot\|_{F}$), we have
$\|\mathrm{diag}(\tilde{\sigma}_{1}-\sigma_{1},\dots,\tilde{\sigma}_{l}-\sigma_{l})\|\leq\|\tilde{\bm{X}}-\bm{X}\|.$
Lemma 12.
Suppose that $(\bm{\sigma},\bm{\Pi})\in\mathbb{R}^{d_{\min}}\times\mathcal{P}^{d_{\min}}$ satisfies the following conditions:
$\begin{cases}\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{d_{\min}}\geq 0,\\ \sigma_{i}^{2L-1}-\sqrt{\lambda}\,y_{\pi(i)}\,\sigma_{i}^{L-1}+\lambda\,\sigma_{i}=0,\ \forall i\in[d_{\min}],\end{cases}$ (191)
where $\pi:[d_{\min}]\to[d_{\min}]$ is the permutation corresponding to $\bm{\Pi}$. Then, the pair $(\bm{\sigma},\bm{\Pi})$ belongs to the set $\mathcal{B}$.
Proof.
According to (191), we conclude
$\sigma_{\pi^{-1}(i)}^{2L-1}-\sqrt{\lambda}\,y_{i}\,\sigma_{\pi^{-1}(i)}^{L-1}+\lambda\,\sigma_{\pi^{-1}(i)}=0,\ \sigma_{\pi^{-1}(i)}\geq 0,\ \forall i\in[d_{\min}].$
Combining this with the definition of $\mathcal{A}$ in (33), we have $\bm{a}:=\bm{\Pi}^{T}\bm{\sigma}=\left(\sigma_{\pi^{-1}(1)},\ \cdots,\ \sigma_{\pi^{-1}(d_{\min})}\right)^{T}\in\mathcal{A}$. Since $\bm{\Pi}\bm{a}=\bm{\sigma}$ and by the definition of $\mathcal{B}$ in (34), it follows that the pair $(\bm{\sigma},\bm{\Pi})\in\mathcal{B}$. ∎
D.1 Proof of (160)
To simplify the notation, for each $i\in[p]$, we introduce the following definition: $\hat{\bm{T}}_{i}:=\prod_{l=2}^{L}\bm{T}_{l}^{(i)}$. Then we have
$\quad\left\|\left(\hat{\bm{\Sigma}}_{1}\hat{\bm{\Sigma}}_{1}^{T}\right)^{L-1}\hat{\bm{\Sigma}}_{1}+\lambda\hat{\bm{\Sigma}}_{1}-\sqrt{\lambda}\,\mathrm{BlkDiag}\left((\bm{\Sigma}_{1}^{(1)})^{L-1},\dots,(\bm{\Sigma}_{1}^{(p)})^{L-1},\bm{0}\right)\bm{\Psi}\hat{\bm{T}}\right\|_{F}$
$=\left\|\mathrm{BlkDiag}\left(\hat{\bm{T}}_{1},\dots,\hat{\bm{T}}_{p},\bm{I}_{d_{1}-r_{\sigma}}\right)\left(\left(\hat{\bm{\Sigma}}_{1}\hat{\bm{\Sigma}}_{1}^{T}\right)^{L-1}\hat{\bm{\Sigma}}_{1}+\lambda\hat{\bm{\Sigma}}_{1}\right)\hat{\bm{T}}^{T}\right.$
$\quad\left.-\sqrt{\lambda}\,\mathrm{BlkDiag}\left(\bm{A}_{1},\dots,\bm{A}_{p},\bm{0}\right)\bm{\Psi}\right\|_{F}$
$\leq\left\|\left(\hat{\bm{\Sigma}}_{1}\hat{\bm{\Sigma}}_{1}^{T}\right)^{L-1}\hat{\bm{\Sigma}}_{1}+\lambda\hat{\bm{\Sigma}}_{1}-\sqrt{\lambda}\,\mathrm{BlkDiag}\left(\bm{A}_{1},\dots,\bm{A}_{p},\bm{0}\right)\bm{\Psi}\right\|_{F}$
$\quad+\sum_{i=1}^{p}\left\|\hat{\bm{T}}_{i}(\bm{\Sigma}_{1}^{(i)})^{2L-1}\hat{\bm{T}}_{i}^{T}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\right\|_{F}+\lambda\sum_{i=1}^{p}\left\|\hat{\bm{T}}_{i}\bm{\Sigma}_{1}^{(i)}\hat{\bm{T}}_{i}^{T}-\bm{\Sigma}_{1}^{(i)}\right\|_{F}$
$\leq\eta_{3}\left\|\nabla G(\bm{W})\right\|_{F}+\sum_{i=1}^{p}\left\|\hat{\bm{T}}_{i}(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\hat{\bm{T}}_{i}\right\|_{F}+\lambda\sum_{i=1}^{p}\left\|\hat{\bm{T}}_{i}\bm{\Sigma}_{1}^{(i)}-\bm{\Sigma}_{1}^{(i)}\hat{\bm{T}}_{i}\right\|_{F}$
$\leq\left(\eta_{3}+\frac{p\eta_{1}(2L-1)L}{\sigma^{*}_{\min}}\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{2L-2}+\frac{\lambda p\eta_{1}L}{\sigma^{*}_{\min}}\right)\left\|\nabla G(\bm{W})\right\|_{F},$
where the first equality follows from the definition of $\bm{A}_{i}$ for each $i\in[p]$ in (71); the second inequality follows from (159); and the last inequality follows from (134) and the following fact(see (192) and (193)). For each i in $[p]$, we have:
$\quad\left\|\hat{\bm{T}}_{i}(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\hat{\bm{T}}_{i}\right\|_{F}=\left\|\left(\prod_{l=2}^{L}\bm{T}_{l}^{(i)}\right)(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\prod_{l=2}^{L}\bm{T}_{l}^{(i)}\right\|_{F}$ (192)
$\leq\left\|\left(\prod_{l=2}^{L-1}\bm{T}_{l}^{(i)}\right)\left(\bm{T}_{L}^{(i)}(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\bm{T}_{L}^{(i)}\right)\right\|_{F}+$
$\quad\left\|\left(\prod_{l=2}^{L-2}\bm{T}_{l}^{(i)}\right)\left(\bm{T}_{L-1}^{(i)}(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\bm{T}_{L-1}^{(i)}\right)\bm{T}_{L}^{(i)}\right\|_{F}+\cdots$
$\quad+\left\|\left(\bm{T}_{2}^{(i)}(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\bm{T}_{2}^{(i)}\right)\left(\prod_{l=3}^{L}\bm{T}_{l}^{(i)}\right)\right\|_{F}$
$\leq\sum_{l=2}^{L}\sum_{j=1}^{2L-1}\left\|(\bm{\Sigma}_{1}^{(i)})^{j-1}\left(\bm{T}_{l}^{(i)}\bm{\Sigma}_{1}^{(i)}-\bm{\Sigma}_{1}^{(i)}\bm{T}_{l}^{(i)}\right)(\bm{\Sigma}_{1}^{(i)})^{2L-j-1}\right\|_{F}$
$\leq\frac{\eta_{1}(2L-1)L}{\sigma^{*}_{\min}}\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{2L-2}\|\nabla G(\bm{W})\|_{F},$
where the last inequality follows from (134) and (57). Similarly, we have

### v3

@@ B.3.1 Proof of (97) and (98)
We first bound
$\left\|(\hat{\bm{\Sigma}}_{L}\hat{\bm{\Sigma}}_{L}^{T})^{L-1}\hat{\bm{\Sigma}}_{L}+\lambda\hat{\bm{\Sigma}}_{L}-\sqrt{\lambda}\bm{\Psi}\mathrm{BlkD}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{0}\right)\right\|_{F}$
$\leq$ $\left\|(\bm{\Sigma}_{L}\bm{\Sigma}_{L}^{T})^{L-1}\bm{\Sigma}_{L}+\lambda\bm{\Sigma}_{L}-\sqrt{\lambda}\bm{\Psi}\mathrm{BlkD}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{B}_{p+1}\right)\right\|_{F}+\lambda\|\hat{\bm{\Sigma}}_{L}-\bm{\Sigma}_{L}\|_{F}$
$\leq$ $\left(c_{2}+\left(\left(\frac{3\sigma_{\max}^{*}}{2}\right)^{2(L-1)}+\lambda\right)c_{3}\sqrt{d_{\max}}\right)\left\|\nabla G(\bm{W})\right\|_{F}+\sqrt{\lambda}y_{1}\|\bm{B}_{p+1}\|_{F}$ (136)
$\leq$ $\eta_{3}\left\|\nabla G(\bm{W})\right\|_{F},$ (137)
where the second inequality follows from (55), (59), (134), and (135), and the last uses (35) and (55). Using (55) and (140), we have
$\leq$ $\left\|(\hat{\bm{\Sigma}}_{L}\hat{\bm{\Sigma}}_{L}^{T})^{L-1}\hat{\bm{\Sigma}}_{L}+\lambda\hat{\bm{\Sigma}}_{L}-\sqrt{\lambda}\bm{\Psi}\hat{\bm{T}}\mathrm{BlkD}\left((\bm{\Sigma}_{L}^{(1)})^{L-1},\dots,(\bm{\Sigma}_{L}^{(p)})^{L-1},\bm{0}\right)\right\|_{F}^{2}$
$=$ $\left\|(\hat{\bm{\Sigma}}_{L}\hat{\bm{\Sigma}}_{L}^{T})^{L-1}\hat{\bm{\Sigma}}_{L}+\lambda\hat{\bm{\Sigma}}_{L}-\sqrt{\lambda}\bm{\Psi}\mathrm{BlkD}\left(\bm{B}_{1},\dots,\bm{B}_{p},\bm{0}\right)\right\|_{F}^{2}$
Dividing the above inequality by $\sqrt{\lambda}\,\sigma_{r_{\sigma}}^{L-1}(\bm{W}_{L})$ and using (137) and the auxiliary function $\varphi$ in (85) yields
[…]
which is just the desired inequality (97). By applying the same argument used to derive (137) using (58), we have
$\left\|(\hat{\bm{\Sigma}}_{1}\hat{\bm{\Sigma}}_{1}^{T})^{L-1}\hat{\bm{\Sigma}}_{1}+\lambda\hat{\bm{\Sigma}}_{1}-\sqrt{\lambda}\mathrm{BlkD}\left(\bm{A}_{1},\dots,\bm{A}_{p},\bm{0}\right)\bm{\Psi}\right\|_{F}\leq\eta_{3}\left\|\nabla G(\bm{W})\right\|_{F}.$ (139)
Now, we claim
$\left\|\left(\hat{\bm{\Sigma}}_{1}\hat{\bm{\Sigma}}_{1}^{T}\right)^{L-1}\hat{\bm{\Sigma}}_{1}+\lambda\hat{\bm{\Sigma}}_{1}-\sqrt{\lambda}\,\mathrm{BlkD}\left((\bm{\Sigma}_{1}^{(1)})^{L-1},\dots,(\bm{\Sigma}_{1}^{(p)})^{L-1},\bm{0}\right)\bm{\Psi}\hat{\bm{T}}\right\|_{F}\leq\eta_{4}\left\|\nabla G(\bm{W})\right\|_{F}.$ (140)
Applying a similar argument as in (138) to the rows of the above inequality yields (98).
In the rest of this proof, we show that (140) holds. For each $i\in[p]$, we define $\hat{\bm{T}}_{i}:=\prod_{l=2}^{L}\bm{T}_{l}^{(i)}$. For each $i\in[p]$, we have
$\textstyle\quad\left\|\hat{\bm{T}}_{i}(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\hat{\bm{T}}_{i}\right\|_{F}$
$=\left\|\left(\prod_{l=2}^{L}\bm{T}_{l}^{(i)}\right)(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\prod_{l=2}^{L}\bm{T}_{l}^{(i)}\right\|_{F}$
$\textstyle\leq\left\|\left(\prod_{l=2}^{L-1}\bm{T}_{l}^{(i)}\right)\left(\bm{T}_{L}^{(i)}(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\bm{T}_{L}^{(i)}\right)\right\|_{F}$ (141)
$\quad\textstyle+\left\|\left(\prod_{l=2}^{L-2}\bm{T}_{l}^{(i)}\right)\left(\bm{T}_{L-1}^{(i)}(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\bm{T}_{L-1}^{(i)}\right)\bm{T}_{L}^{(i)}\right\|_{F}+\cdots$
$\textstyle\quad+\left\|\left(\bm{T}_{2}^{(i)}(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\bm{T}_{2}^{(i)}\right)\left(\prod_{l=3}^{L}\bm{T}_{l}^{(i)}\right)\right\|_{F}$
$\textstyle\leq\sum_{l=2}^{L}\sum_{j=1}^{2L-1}\left\|(\bm{\Sigma}_{1}^{(i)})^{j-1}\left(\bm{T}_{l}^{(i)}\bm{\Sigma}_{1}^{(i)}-\bm{\Sigma}_{1}^{(i)}\bm{T}_{l}^{(i)}\right)(\bm{\Sigma}_{1}^{(i)})^{2L-j-1}\right\|_{F}$
$\textstyle\leq\frac{\eta_{1}(2L-1)L}{\sigma^{*}_{\min}}\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{2L-2}\|\nabla G(\bm{W})\|_{F},$ (142)
where the last inequality follows from (35) and (65). Similarly, we have
Then we have
$\textstyle\quad\left\|\left(\hat{\bm{\Sigma}}_{1}\hat{\bm{\Sigma}}_{1}^{T}\right)^{L-1}\hat{\bm{\Sigma}}_{1}+\lambda\hat{\bm{\Sigma}}_{1}-\sqrt{\lambda}\,\mathrm{BlkD}\left((\bm{\Sigma}_{1}^{(1)})^{L-1},\dots,(\bm{\Sigma}_{1}^{(p)})^{L-1},\bm{0}\right)\bm{\Psi}\hat{\bm{T}}\right\|_{F}$
$\textstyle=\left\|\mathrm{BlkD}\left(\hat{\bm{T}}_{1},\dots,\hat{\bm{T}}_{p},\bm{I}_{d_{1}-r_{\sigma}}\right)\left(\left(\hat{\bm{\Sigma}}_{1}\hat{\bm{\Sigma}}_{1}^{T}\right)^{L-1}\hat{\bm{\Sigma}}_{1}+\lambda\hat{\bm{\Sigma}}_{1}\right)\hat{\bm{T}}^{T}-\sqrt{\lambda}\,\mathrm{BlkD}\left(\bm{A}_{1},\dots,\bm{A}_{p},\bm{0}\right)\bm{\Psi}\right\|_{F}$
$\textstyle\leq\left\|\left(\hat{\bm{\Sigma}}_{1}\hat{\bm{\Sigma}}_{1}^{T}\right)^{L-1}\hat{\bm{\Sigma}}_{1}+\lambda\hat{\bm{\Sigma}}_{1}-\sqrt{\lambda}\,\mathrm{BlkD}\left(\bm{A}_{1},\dots,\bm{A}_{p},\bm{0}\right)\bm{\Psi}\right\|_{F}$
$\textstyle\quad+\sum_{i=1}^{p}\left\|\hat{\bm{T}}_{i}(\bm{\Sigma}_{1}^{(i)})^{2L-1}\hat{\bm{T}}_{i}^{T}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\right\|_{F}+\lambda\sum_{i=1}^{p}\left\|\hat{\bm{T}}_{i}\bm{\Sigma}_{1}^{(i)}\hat{\bm{T}}_{i}^{T}-\bm{\Sigma}_{1}^{(i)}\right\|_{F}$
$\textstyle\leq\eta_{3}\left\|\nabla G(\bm{W})\right\|_{F}+\sum_{i=1}^{p}\left\|\hat{\bm{T}}_{i}(\bm{\Sigma}_{1}^{(i)})^{2L-1}-(\bm{\Sigma}_{1}^{(i)})^{2L-1}\hat{\bm{T}}_{i}\right\|_{F}+\lambda\sum_{i=1}^{p}\left\|\hat{\bm{T}}_{i}\bm{\Sigma}_{1}^{(i)}-\bm{\Sigma}_{1}^{(i)}\hat{\bm{T}}_{i}\right\|_{F}$
$\textstyle\leq\left(\eta_{3}+\frac{p\eta_{1}(2L-1)L}{\sigma^{*}_{\min}}\left(\frac{3\sigma^{*}_{\max}}{2}\right)^{2L-2}+\frac{\lambda p\eta_{1}L}{\sigma^{*}_{\min}}\right)\left\|\nabla G(\bm{W})\right\|_{F}=\eta_{4}\left\|\nabla G(\bm{W})\right\|_{F},$
where the first equality follows from the definition of $\bm{A}_{i}$ for each $i\in[p]$ in (54), the second inequality follows from (139), and the last inequality follows from (65), (142), and (143). ∎
B.4 Auxiliary Lemma for Proving 3
Lemma B.6.
Let $\bm{\sigma}^{*}\in\mathcal{A}_{\rm sort}\setminus\{\bm{0}\}$ be arbitrary. Then there exists a constant $L_{G}>0$ such that
$\|\nabla G(\bm{W})\|_{F}\;\leq\;L_{G}\,\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}}),\quad\text{for all }\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})\leq\sigma_{\max}^{*},$
where
$L_{G}\;=\;\lambda L\;+\;2^{2L-1}L^{2}(\sigma_{\max}^{*})^{2L-2}\;+\;\sqrt{\lambda}\,y_{1}\,2^{L-2}L^{2}(\sigma_{\max}^{*})^{L-2}.$ (144)
Proof.
@@ B.4 Auxiliary Lemma for Proving 3
Suppose $\|\bm{W}-\bm{W}^{*}\|_{F}=\mathrm{dist}(\bm{W},\mathcal{W}_{\bm{\sigma}^{*}})$ with $\bm{W}^{*}\in\mathcal{W}_{\bm{\sigma}^{*}}$, for each $l\in[L]$, we have
$\|\nabla_{l}G(\bm{W})\|_{F}$
$=$ $\|\nabla_{l}G(\bm{W})-\nabla_{l}G(\bm{W}^{*})\|_{F}$
$\leq$ $\lambda\|\bm{W}_{l}\!-\!\bm{W}^{*}_{l}\|_{F}\!+\!\|\bm{W}_{L:l+1}^{T}(\bm{W}_{L:1}-\sqrt{\lambda}\bm{Y})\bm{W}_{l-1:1}^{T}\!\!-\!\!\bm{W}_{L:l+1}^{*T}(\bm{W}_{L:1}^{*}\!\!-\!\!\sqrt{\lambda}\bm{Y})\bm{W}_{l-1:1}^{*T}\|_{F}$
$\leq$ $\lambda\|\bm{W}_{l}-\bm{W}^{*}_{l}\|_{F}+\|\bm{W}_{L:l+1}^{T}\bm{W}_{L:1}\bm{W}_{l-1:1}^{T}-\bm{W}_{L:l+1}^{*T}\bm{W}_{L:1}^{*}\bm{W}_{l-1:1}^{*T}\|_{F}$
$+\sqrt{\lambda}\|\bm{W}_{L:l+1}^{T}\bm{Y}\bm{W}_{l-1:1}-\bm{W}_{L:l+1}^{*T}\bm{Y}\bm{W}_{l-1:1}^{*T}\|_{F}.$ (145)
By 1, we have $\|\bm{W}_{l}^{*}\|\leq{\sigma}_{\max}^{*}\ \text{and}\ \|\bm{W}_{l}\|\leq\|\bm{W}_{l}^{*}\|+\|\bm{W}-\bm{W}^{*}\|_{F}\leq 2{\sigma}_{\max}^{*},\ \forall l\in[L].$ Using the triangle inequality and the matrix norm inequality $\|\bm{A}\bm{B}\|_{F}\leq\|\bm{A}\|\|\bm{B}\|_{F}$, we have
$\left\|\bm{W}_{L:l+1}^{T}\bm{W}_{L:1}\bm{W}_{l-1:1}^{T}-\bm{W}_{L:l+1}^{*T}\bm{W}_{L:1}^{*}\bm{W}_{l-1:1}^{*T}\right\|_{F}$
$\leq$ $\left\|(\bm{W}_{l+1}-\bm{W}_{l+1}^{*})^{T}\bm{W}_{L:l+2}^{T}\bm{W}_{L:1}\bm{W}_{l-1:1}^{T}\right\|_{F}$
$+\left\|\bm{W}_{l+1}^{*T}(\bm{W}_{l+2}-\bm{W}_{l+2}^{*})^{T}\bm{W}_{L:l+3}^{T}\bm{W}_{l-1:1}^{T}\right\|_{F}+\cdots$
$+\left\|\bm{W}_{L:l+1}^{*T}\bm{W}_{L:1}^{*}\bm{W}_{l-2:1}^{*T}(\bm{W}_{l-1}-\bm{W}_{l-1}^{*})\right\|_{F}$
$\leq$ $L2^{2L-1}{\sigma_{\max}^{*}}^{2L-2}\|\bm{W}-\bm{W}^{*}\|_{F}.$ (146)
Using a similar telescope inequality, we have
$\sqrt{\lambda}\|\bm{W}_{L:l+1}^{T}\bm{Y}\bm{W}_{l-1:1}-\bm{W}_{L:l+1}^{*T}\bm{Y}\bm{W}_{l-1:1}^{*T}\|_{F}\leq\sqrt{\lambda}Ly_{1}2^{L-2}{\sigma}_{\max}^{*L-2}\|\bm{W}-\bm{W}^{*}\|_{F}.$ (147)
Plugging (146) and (147) into (145), we obtain
$\|\nabla G(\bm{W})\|_{F}$ $\leq\sum_{l=1}^{L}\|\nabla G_{l}(\bm{W})\|_{F}$
$\leq\left(\lambda L+2^{2L-1}L^{2}{\sigma}_{\max}^{*2L-2}+\sqrt{\lambda}y_{1}2^{L-2}L^{2}{\sigma}_{\max}^{*L-2}\right)\|\bm{W}-\bm{W}^{*}\|_{F}.$
@@ Appendix C Linear Convergence under the Error Bound
    (Sufficient Decrease) There exists $\kappa_{1}>0$ with
@@
    (Cost-to-Go Estimate) There exists $\kappa_{2}>0$ such that
@@
    (Safeguard) There exists a constant $\kappa_{3}>0$ such that $\|\nabla F(\bm{W}^{k})\|_{F}\leq\kappa_{3}\|\bm{W}^{k+1}-\bm{W}^{k}\|_{F}.$
Then, $\{F(\bm{W}^{k})\}_{k\geq k_{1}}$ converges Q-linearly to $F(\bm{W}^{*})$ and $\{\bm{W}^{k}\}_{k\geq k_{1}}$ converges R-linearly to $\bm{W}^{*}$ for some $\bm{W}^{*}\in\mathcal{W}$.
Most gradient-based algorithms generate sequences that satisfy the above supposition.
@@ Appendix D Auxiliary Results
Lemma D.7.
Let $\bm{A},\hat{\bm{A}}\in\mathbb{R}^{p\times p}$ be symmetric with eigenvalues $\lambda_{1}\geq\cdots\geq\lambda_{p}$ and $\hat{\lambda}_{1}\geq\cdots\geq\hat{\lambda}_{p}$ respectively. Fix $1\leq r\leq s\leq p$ and assume that $\min\{\lambda_{r-1}-\lambda_{r},\lambda_{s}-\lambda_{s+1}\}>0$, where $\lambda_{0}:=\infty$ and $\lambda_{p+1}:=$-$\infty$. Let d:=sr+1d:=s-r+1 and $\bm{V}=(\bm{v}_{r},\bm{v}_{r+1},\dots,\bm{v}_{s})\in\mathbb{R}^{p\times d}$ and $\hat{\bm{V}}=(\hat{\bm{v}}_{r},\hat{\bm{v}}_{r+1},\dots,\hat{\bm{v}}_{s})\in\mathbb{R}^{p\times d}$ have orthonormal columns satisfying $\bm{A}\bm{v}_{j}=\lambda_{j}\bm{v}_{j}$ and $\hat{\bm{A}}\hat{\bm{v}}_{j}=\hat{\lambda}_{j}\hat{\bm{v}}_{j}$ for all $j=r,r+1,\ldots,s$. Then, there exists an orthogonal matrix $\bm{O}\in\mathcal{O}^{d}$ such that
@@
Lemma D.8 (Mirsky’s Inequality [38]).
Let $\bm{X},\tilde{\bm{X}}\in\mathbb{R}^{m\times n}$ be arbitrary matrices with singular values
@@
respectively, where $l=\min\{m,n\}$. Then, for any unitarily invariant norm (e.g., the Frobenius norm), we have
$\left\|\mathrm{diag}(\tilde{\sigma}_{1}-\sigma_{1},\dots,\tilde{\sigma}_{l}-\sigma_{l})\right\|\leq\|\tilde{\bm{X}}-\bm{X}\|.$